I have 60 inches of width and have 48 inches of height in my entertainment center. I am going HDTV shopping, what is the biggest diagonal TV that I can get?
Answer by nerdybill(7384)

I have 60 inches of width and have 48 inches of height in my entertainment center. I am going HDTV shopping, what is the biggest diagonal TV that I can get?
.
That would be the hypotenuse or diagonal:
d^2 = 60^2 + 48^2
d^2 = 3600 + 2304
d^2 = 5904
d = 76.84 inches